# Chain-structured neural architecture search for financial time series forecasting

### v1

@@ Abstract
We compare three popular neural architecture search strategies on chain-structured search spaces: Bayesian optimization, the hyperband method, and reinforcement learning in the context of financial time series forecasting.
@@ 1 Introduction
Deep neural networks have been very successful in a wide variety of tasks over the last two decades. In large part their success is attributed to their ability to perform very well without major manual feature engineering required when compared to more classical techniques. The hierarchical structure of the networks weighs and extracts important features automatically from the data during the learning process [GBC16]. However, the exact architecture of the neural network still has to be prescribed manually by the user. This led to the development of so-called auto-ML techniques that aim to automate this process. In the context of deep neural networks, auto-ML has a very large overlap with neural architecture search (NAS), itself having a large overlap with hyperparameter optimization.
A lot of research has been done in NAS in recent years, see [${}^{+}$23] for an overview and insights from over 1000 papers. However, most work focused on computer vision or natural language applications, with less investigation into architectures for analyzing time series data. In this work, we attempt to bridge this gap by evaluating and comparing the performance of three popular simple NAS strategies on 3 distinct yet similar time series datasets.
A seminal survey article on NAS [EMH19] breaks down the search into 3 components: search spaces, search strategies and performance estimation strategies. Once a search space is set up, an algorithm following the search strategy explores the search space looking for the best neural network configuration, evaluated and eventually selected by the performance estimation strategy. In Section 2, we discuss the exact practical problem from project partner Predictive Layer SA that our neural networks were trying to solve. We discuss the search spaces we set up for NAS in Section 3, the search strategies in Section 5 and the performance estimation together with the general challenges of our problem in Section 4. Our results are summarized and discussed in Section 6 and Section 7 respectively. Finally, we outline possible future directions for this research in Section 8.
@@ 2 Data and problem formulation
The challenge comes with financial markets being notoriously hard to predict [HH19]. Our three source datasets that we report on in Section 6 are for Japanese, German, and US bonds and come with an additional challenge of very high number of features (around 1000 each), comparable with the number of training examples (under 4000), which is very low in a deep learning setting [GBC16], especially relatively to the number of features. To circumvent the high dimensionality, the number of features was reduced approximately by a factor of $3$ by removing features in the prepared datasets ‘time-derived’ from original features, such as by lagging original features by fixed time steps or taking the mean over the same time steps. Such time-derived features are standard in the industry [HH19], but are fundamentally used because the algorithms applied to the data can only take in one row of data as input. All the more sophisticated neural network architectures take in a sequence of rows of data, making time-derived features redundant. See Section 3 for more details. Additionally, principal component analysis (PCA) was applied to further reduce the number of features. Based on the scree plot coming from PCA analysis, two breakpoints were tested: the point of highest curvature (the usual choice) and the point where the graph flattens, retaining most of the information. Thus for each of the 3 source datasets, we obtained 4 variations: no modification, time-derived features removed and 2 PCA transformations, resulting in 12 datasets total.
@@ 3 Architecture types and their search spaces
In NAS, the search spaces and search strategies go hand in hand and are designed together essentially at the same time. The simplest among the search spaces are chain-structured. Search spaces with a chain structure feature a straightforward architectural topology: a sequential chain of operational layers. These configurations frequently utilize cutting-edge manually designed architectures as their foundational framework, trying to find the best configuration for the specific dataset at hand by essentially hyperparameter optimization. Despite their simplicity, chain-structured search spaces often yield very good results [${}^{+}$23]. The biggest downsides of this approach when compared to the more complicated cell-based and one-shot approaches are its limited flexibility and scalability when applied to very large datasets. While being limited to a pre-defined overall structure prevents discovering truly novel architectures, in practice well-designed and well-understood architectures usually perform best [${}^{+}$23]. Scalability becomes important when dealing with large image and natural language datasets that most research on NAS focuses on, however in the case of time series and especially our small datasets, our neural networks train in seconds. Therefore scalability was not an issue for us, and thus we focused on chain-structured search spaces.
For the foundational frameworks the search spaces were built upon, we concentrated on simple well-known architecture types, like feedforward (FFNN, also known as multi-layer perceptron), convolutional (CNN), and recurrent (RNN) networks. This was done both for their proven perforamnce in time-series context and due to them having fewer parameters than most cutting-edge architectures, which is more appropriate when dealing with very little data as in our case. However, we also examined the state of the art Temporal Fusion Transformer (TFT) architecture for comparison (see below). Each of these architectures was adapted to the problem at hand and for each of them a search space of hyperparameters was set up to determine the exact configuration. The search strategies discussed in Section 5 then optimized each of these separately, and the best-performing ones were compared. This approach is similar to what [ZL17] did with a single type of CNN for image classification.
@@ 3.1 Feedforward networks
Feedforward networks have the advantage of being the simplest type of deep neural networks. The main disadvantage for our problem comes from the fact that they cannot take the time-dimension into account explicitly. Feedforward networks take a row of numbers as input, one row of our datasets corresponding to one day information for various time series. Implicitly, the time information is taken into account in the form of delayed (or averaged) features – a column in the dataset that is essentially the same as another column with an added delay of a couple days (or averaged over a couple days, respectively), a standard practice in the field [HH19].
@@ 3.2 Convolutional neural networks
Although designed for and used mainly with images, convolutional neural networks (CNNs) are suitable for any kind of grid-like data, especially when there is a natural fixed notion of distance between neighboring data points, such as the case for time series [GBC16].
@@
Now, a ‘chunk’ of our dataset – a subtable of fixed length (in time) can be passed as a single input to our model. Evaluating the history from the last, say, $15$ days explicitly, the model is making a prediction for the following day as before.
@@
Both one- and two-dimensional CNNs can be used for our problem. One-dimensional (1D) CNNs only perform the convolution operation over the time dimension, treating separate features as parallel input channels, just like the red, green and blue channels are considered when working with images. Two-dimensional (2D) CNNs take each chunk as a single-channel ‘image’, performing the convolution operation over both the time dimension and across the various features. The latter seems counterintuitive as there is no fixed distance between the features, but [HH19] found success with this technique. After experimenting with 2D CNNs on our data, we found them a lot more computationally expensive than others, while often performing worse. For this reason 2D CNNs were later dropped from our consideration.
@@
The length in time of the chunks passed as individual examples (‘chunk length’) was a crucial hyperparameter to optimize for all of the CNNs. The number of convolutional layers (including pooling and activation functions) was fixed at $3$ for all CNN architectures to reduce the size of the search spaces, but the kernel sizes were optimized. The number of convolution filters per layer was optimized too, but it works differently for different implementations, as discussed below. Dropout layers were used once again for regularization and the dropout rate was optimized too, together with the learning rate.
3.2.1 Depthwise 1D CNNs
@@
Among 1D CNNs, there are two distinct ways to perform the convolution operation. In depthwise convolution, each channel (input feature) is passed separately through its own convolutional and pooling layers. Each of these then gives a fixed number (the number of separate convolution filters) of output channels per input channel, interpreted as a summary of what happened to this feature in the time span considered. The convolution and pooling operations will shrink the length (in time) of each output channel, possibly to $1$ (otherwise flattening is applied), and all of these are then passed to a single dense layer that makes a prediction based on these summaries.
Although easy to interpret, this method had the problem of blowing up the dimensionality of the problem, as each of the features gets its own set of convolution parameters. This made computation infeasible and more work is needed to be carried out for this architecture.
3.2.2 Ordinary 1D CNNs
In basic (we call them ‘ordinary’) 1D convolution, all input channels are mixed already in the first (and subsequent) convolution layers. The output of convolution is flattened and passed to a single dense layer for making the final prediction.
@@ 3.3 Recurrent neural networks
Recurrent neural networks (RNNs) were designed for and are very successful at treating sequential data [GBC16] and are thus the natural choice for our problem. Both simple RNNs and more sophisticated long short-term memory (LSTM) RNNs were implemented and optimized separately, but they have the same search spaces of hyperparameters. In both cases, stacked RNNs were used, where the output of one RNN is taken as input to the next one. The number of such stacked layers was a hyperparameter to optimize. Each RNN layer also has its number of hidden units, which was also optimized. The chunk length, here naturally interpreted as the length of the input sequence, was once again optimized, together with dropout and learning rates as before.
LSTMs provide many advantages over basic RNNs [GBC16] without too much additional computational cost, especially in the case of our small datasets. For this reason after initial trials, we proceeded with just LSTMs for the final tests.
@@ 3.4 Temporal fusion transformer
Forecasting across multiple time horizons often involves a intricate amalgamation of inputs, encompassing static (i.e., time-invariant) covariates, known future inputs, and additional exogenous time series observed solely in the past. The challenge lies in handling this complexity without prior knowledge of how these inputs interact with the target variable. While various deep learning approaches have been proposed, they often manifest as ‘black-box’ models, lacking transparency regarding their utilization of the diverse input types encountered in practical scenarios.
In a recent work [LALP21], the Temporal Fusion Transformer (TFT) was introduced as an innovative attention-based architecture addressing this issue. TFT not only achieves high-performance multi-horizon forecasting but also provides interpretable insights into temporal dynamics. In order to capture temporal relationships at different scales, TFT incorporates recurrent layers for local processing and interpretable self-attention layers for modelling long-term dependencies. Specialized components within TFT are employed to select relevant features, and a series of gating layers effectively suppress unnecessary components, resulting in high performance across a broad spectrum of scenarios.
[LALP21] demonstrated significant performance enhancements over existing benchmarks across various real-world datasets. The study also highlighted three practical use cases illustrating the interpretability of TFT, showcasing its efficacy in shedding light on the decision-making process. The PyTorch Forecasting package [PyT23] has on open-source implementation of the model. It comes with its own optimizer, selecting the optimal number of attention heads, the network’s hidden size, and learning and dropout rates.
@@
4 Challenges and performance estimation
There are many evaluation metrics applicable to the binary classification problem. Although common for other classification tasks, accuracy ($ACC$) is not a good metric as a trivial predictor that always gives the same output no matter the input can have high accuracy in the case of an unbalanced dataset. Balanced accuracy ($bACC$), the average between the true negative and true positive rates, mitigates this problem. Another popular alternative is the $F_{1}$ score, the harmonic mean of precision and recall. The area under the receiver operating characteristic (ROC) curve, or $AUC$ for short is another good choice. It is hard to say which one is better, and a good strategy could mix a combination of these.
Predicting financial markets is extremely challenging and this is reflected in our test metrics always being very close to $0.5$, i.e. the predictors being very close to random. This would be regarded as bad performance in many areas, but for financial datasets it is quite the norm: many companies make successful investments using algorithms (whether traditional or ML-based) being only marginally better than a random predictor [HH19]. Further, neural networks in particular in this context suffer from comparatively high variance depending on the random seed chosen during training due to their stochastic nature [HH19]. Even identical neural network architectures, trained in exactly the same way but with different random seeds can have vastly different performance on the test dataset, e.g. $0.56$ on a ‘good’ random seed and $0.48$ (i.e. worse than random) on a ‘bad’ one. This is a problem for NAS as what seemed like a good (or bad) neural network configuration might have just been a lucky (or unlucky respectively) seed. To mitigate this problem, we trained and tested each architecture configuration chosen $15$ times with different random seeds, averaging the test metrics and using the averages as our guide. At the end of optimization, the best performing configuration was trained and tested $50$ times (again, with different random seeds every time) to further eliminate random seed variation.
@@ 5 Search strategies
We started with $7$ different broad architecture types: FFNN, depthwise 1D CNN, ordinary 1D CNN, 2D CNN, simple RNN, LSTM, and TFT, but for final optimization studies only FFNN, ordinary 1D CNN, and LSTM were used, as discussed in Section 3. Each of them have their own search space of hyperparameters determining the exact network architecture.
Classical approaches for hyperparameter tuning in machine learning, such as grid search and random search, have been widely used due to their simplicity [BB12]. However, these methods can be computationally expensive and inefficient, especially when dealing with high-dimensional search spaces or expensive objective functions [BBBK11]. The survey on NAS [${}^{+}$23] favourably featured three other popular search strategies: Bayesian optimization, reinforcement learning, and the hyperband method which are the three methods we explored extensively.
Fundamentally, all of these suggest a specific configuration of hyperparameters of a neural network. The network is then trained and tested (in our case repeatedly so, see Section 4), the results saved, and the algorithm suggests the next configuration to try. This continues for a set number of trials (in our studies, we set it at $300$) and in the end the best performing configuration is selected.
@@ 5.1 Bayesian optimization
Bayesian optimization has emerged as a popular alternative to grid and random search for hyperparameter tuning, as it efficiently explores the search space and intelligently guides the optimization process [SLA12]. One widely used technique in Bayesian optimization is based on Gaussian Processes (GP), which fit the objective function using Gaussian processes as probabilistic models and leverages acquisition functions to balance exploration and exploitation [RW06]. GPs provide a flexible, non-parametric tool for modeling complex functions, allowing for uncertainty quantification and adaptation to new data [${}^{+}$16].
However, GP-based methods have some limitations, such as the inability to handle categorical features or dependent parameters without ad hoc modifications [BCdF10]. This has led to the development of alternative methods, such as the Tree-structured Parzen Estimator (TPE), which has gained attention for its scalability and effectiveness [BBBK11]. TPE models the search space using hierarchical Parzen estimators, which adaptively partition the space to model complex, high-dimensional functions [BBBK11].
TPE is particularly suitable for addressing the shortcomings of GP-based methods, as it can naturally handle categorical features and dependent parameters without requiring extensive adaptations. Performance comparisons between TPE, GP-based methods, and classical search techniques have shown that TPE often yields improved results across various tasks and datasets [${}^{+}$13].
A notable implementation of TPE is available in the Optuna package, a versatile optimization library that has demonstrated its utility in various machine learning tasks [${}^{+}$19].

### v2

@@ Abstract
Neural architecture search (NAS) emerged as a way to automatically optimize neural networks for a specific task and dataset. Despite an abundance of research on NAS for images and natural language applications, similar studies for time series data are lacking. Among NAS search spaces, chain-structured are the simplest and most applicable to small datasets like time series. We compare three popular NAS strategies on chain-structured search spaces: Bayesian optimization (specifically Tree-structured Parzen Estimator), the hyperband method, and reinforcement learning in the context of financial time series forecasting. These strategies were employed to optimize simple well-understood neural architectures like the MLP, 1D CNN, and RNN, with more complex temporal fusion transformers (TFT) and their own optimizers included for comparison. We find Bayesian optimization and the hyperband method performing best among the strategies, and RNN and 1D CNN best among the architectures, but all methods were very close to each other with a high variance due to the difficulty of working with financial datasets. We discuss our approach to overcome the variance and provide implementation recommendations for future users and researchers.
Keywords: neural architecture search, time series forecasting, hyperparameter optimization, deep learning, neural networks, reinforcement learning.
This is the accepted version of the paper published in International Journal of Data Science and Analytics.
DOI: 10.1007/s41060-024-00690-y
@@ 1 Introduction
Deep neural networks have been very successful in a wide variety of tasks over the last two decades. In large part their success is attributed to their ability to perform very well without major manual feature engineering required when compared to more classical techniques. [1]. However, the exact architecture of the neural network still has to be prescribed manually by the user. This led to the development of so-called auto-ML techniques that aim to automate this process. In the context of deep neural networks, auto-ML has a very large overlap with neural architecture search (NAS), itself having a large overlap with hyperparameter optimization.
A lot of research has been done in NAS in recent years, see [2] for an overview and insights from over 1000 papers. However, most work focused on computer vision or natural language applications, with less investigation into architectures for analyzing time series data. In this work, we attempt to bridge this gap by evaluating and comparing the performance of three popular simple NAS strategies on 3 distinct yet similar (financial) time series datasets.
Modern competitive neural nets perform very well on natural language and images datasets, and any improvement is marginal, while the datasets and networks are huge, requiring vast resources to train each individual network and thus directing NAS research toward more complicated methods that reduce the number of overall trials [2, 3]. Meanwhile, time series, and especially financial time series,x pose a fundamentally different challenge. With time series, the datasets are often small due to the low number of historical samples, which forces the networks to have fewer parameters too in order to avoid overfitting. This results in neural networks that train in seconds on modern hardware. The performance, especially in the case of financial time series, on the other hand, is always comparatively very poor. The challenge comes with financial markets being notoriously hard to predict. Intuitively, this is clear: if it was easy, every researcher would have been a millionaire. The problem is challenging due to the inherently noisy and non-stationary nature of financial time series. Market prices are influenced by a myriad of unpredictable factors, such as economic events, investor behavior, and geopolitical developments, leading to a low signal-to-noise ratio. Consequently, even state-of-the-art algorithms struggle to achieve high performance, typically resulting in F1, (balanced) accuracy, and AUC scores well below 0.6 on average when applied to test data [4, 5, 6].
A seminal survey article on NAS [3] breaks down the search into 3 components: search spaces, search strategies, and performance estimation strategies. Once a search space is set up, an algorithm following the search strategy explores the search space looking for the best neural network configuration, evaluated and eventually selected by the performance estimation strategy. Following this terminology, we discuss the search spaces we set up for NAS in Section 3, the search strategies in Section 4 and the performance estimation together with methodology in Section 5. In Section 2 we discuss the exact practical problem from project partner Predictive Layer SA that our neural networks were trying to solve. Our results are summarized and discussed in Section 6 and Section 7, respectively. Finally, we outline possible future directions for this research in Section 7.1.
@@ 2 Data and problem formulation
Our three source datasets that we report on in Section 6 are for Japanese, German, and US bonds. Each of the tabular source datasets has a size of about $4000\times 1000$: $1000$ columns (input features), representing the many financial markers believed to be predictive of the target, with each row corresponding to a business day over a 15-year period. While this sample size is relatively small for deep learning applications [1] – especially given the high number of features – such sizes are typical in time series datasets. Thus we hope that our findings will be relevant to time series practitioners.
@@ 3 Architecture types and their search spaces
In NAS, the search spaces and search strategies go hand in hand and are designed together essentially at the same time. The simplest among the search spaces are chain-structured. Search spaces with a chain structure feature a straightforward architectural topology: a sequential chain of operational layers. These configurations frequently utilize cutting-edge manually designed architectures as their foundational framework, trying to find the best configuration for the specific dataset at hand by essentially hyperparameter optimization. Despite their simplicity, chain-structured search spaces often yield very good results [2]. The biggest downsides of this approach when compared to the more complicated cell-based and one-shot approaches are its limited flexibility and scalability when applied to very large datasets. While being limited to a pre-defined overall structure prevents discovering truly novel architectures, in practice well-designed and well-understood architectures usually perform best [2]. Scalability becomes important when dealing with large image and natural language datasets that most research on NAS focuses on; however, in the case of time series and especially our small datasets, our neural networks train in seconds. Therefore scalability was not an issue for us, and thus we focused on chain-structured search spaces.
For the foundational frameworks that the search spaces were built upon, we concentrated on simple well-known architecture types, like feedforward (FFNN, also known as multi-layer perceptron or MLP), convolutional (CNN), and recurrent (RNN) networks. This was done both for their proven performance in time-series context and due to them having fewer parameters than most cutting-edge architectures, which is more appropriate when dealing with very little data as in our case. However, we also examined the state of the art Temporal Fusion Transformer (TFT) architecture for comparison (see below). Each of these architectures was adapted to the problem at hand and for each of them a search space of hyperparameters was set up to determine the exact configuration. The search strategies discussed in Section 4 then optimized each of these separately, and the best-performing ones were compared. This approach is similar to what [7] did with a single type of CNN for image classification.
@@ 3.1 Feedforward networks
Feedforward networks (aka multi-layer perceptron or MLP) have the advantage of being the simplest type of deep neural networks. The main disadvantage for our problem comes from the fact that they cannot take the time dimension into account explicitly. Feedforward networks take a row of numbers as input, one row of our datasets corresponding to one day information for various time series. Implicitly, the time information is taken into account in the form of delayed (or averaged) features – a column in the dataset that is essentially the same as another column with an added delay of a couple days (or averaged over a couple days, respectively), a standard practice in the field [5].
@@ 3.2 Convolutional neural networks
Although designed for and used mainly with images, convolutional neural networks (CNNs) are suitable for any kind of grid-like data, especially when there is a natural fixed notion of distance between neighboring data points, such as the case for time series [1].
@@
CNNs allow taking the time-dimension into account explicitly and naturally. Instead of each input being a single row of data corresponding to a single business day (as was the case for FFNN), a sliding window or a ‘chunk’ of our dataset – a subtable comprised of several consecutive rows can be passed as a single input to a CNN. Evaluating the history from the last, say, $15$ business days explicitly, the model is making a prediction for the following day as before.
@@
Both one- and two-dimensional CNNs can be used for our problem. One-dimensional (1D) CNNs only perform the convolution operation over the time dimension, treating separate features as parallel input channels, just like the red, green, and blue channels are considered when working with images. Thus each chunk of data is passed as a 1D ‘image’ of size chunk_length (a hyperparameter to be optimized) and $n_{features}$ (fixed number coming from the dataset) channels. Two-dimensional (2D) CNNs take each chunk as a single-channel 2D ‘image’ of size chunk_length $\times$ $n_{features}$, performing the convolution operation over both the time dimension and across the various features. The latter seems counterintuitive as there is no fixed distance between the features, but such approach was found to be successful in the context of financial time series before[5]. After experimenting with 2D CNNs on our data, we found them a lot more computationally expensive than others, while often performing worse. For this reason 2D CNNs were later dropped from our consideration.
@@
The length in time of the chunks passed as individual examples, chunk_length, was a crucial hyperparameter to optimize for all of the CNNs. The number of convolutional layers (including pooling and activation functions) was fixed at $3$ for all CNN architectures to reduce the size of the search spaces, but the kernel sizes were optimized. The number of convolution filters per layer was optimized too, but it works differently for different implementations, as discussed below. Dropout layers were used once again for regularization and the dropout rate was optimized too, together with the learning rate.
@@
Among 1D CNNs, there are two distinct ways to perform the convolution operation. In depthwise convolution, each channel (input feature) is passed separately through its own convolutional and pooling layers. Each of these then gives a fixed number (the number of separate convolution filters) of output channels per input channel, interpreted as a summary of what happened to this feature in the time span considered. The convolution and pooling operations will shrink the length (in time) of each output channel, possibly to $1$ (otherwise flattening is applied), and all of these are then passed to a fully connected layer that makes a prediction based on these summaries. Although easy to interpret, this method had the problem of blowing up the dimensionality of the problem, as each of the features gets its own set of convolution parameters. This made computation infeasible and we abandoned this approach. In basic (we call them ‘ordinary’) 1D convolution, all input channels are mixed already in the first (and subsequent) convolution layers. The output of convolution is flattened and passed to a single fully connected layer for making the final prediction.
@@ 3.3 Recurrent neural networks
Recurrent neural networks (RNNs) were designed for and are very successful at treating sequential data [1] and are thus the natural choice for our problem. The same chunks of data as for CNNs were passed as individual inputs to the RNNs, here naturally interpreted as sequences of length chunk_length and dimension n_features. Both simple RNNs and more sophisticated long short-term memory (LSTM) RNNs were implemented and optimized separately, but they have the same search spaces of hyperparameters. In both cases, stacked RNNs were used, where the output of one RNN is taken as input to the next one. The number of such stacked layers was a hyperparameter to optimize. Each RNN layer also has its number of hidden units, which was also optimized. The chunk length was once again optimized, together with dropout and learning rates as before.
LSTMs provide many advantages over basic RNNs [1] without too much additional computational cost, especially in the case of our small datasets. For this reason after initial trials, we proceeded with just LSTMs for the final tests.
@@ 3.4 Temporal fusion transformer
Forecasting across multiple time horizons often involves an intricate amalgamation of inputs, encompassing static (i.e., time-invariant) covariates, known future inputs, and additional exogenous time series observed solely in the past. The challenge lies in handling this complexity without prior knowledge of how these inputs interact with the target variable. While various deep learning approaches have been proposed, they often manifest as ‘black-box’ models, lacking transparency regarding their utilization of the diverse input types encountered in practical scenarios.
In a recent work [8], the Temporal Fusion Transformer (TFT) was introduced as an innovative attention-based architecture addressing this issue. TFT not only achieves high-performance multi-horizon forecasting but also provides interpretable insights into temporal dynamics. In order to capture temporal relationships at different scales, TFT incorporates recurrent layers for local processing and interpretable self-attention layers for modeling long-term dependencies. Specialized components within TFT are employed to select relevant features, and a series of gating layers effectively suppress unnecessary components, resulting in high performance across a broad spectrum of scenarios.
[8] demonstrated significant performance enhancements over existing benchmarks across various real-world datasets. The study also highlighted three practical use cases illustrating the interpretability of TFT, showcasing its efficacy in shedding light on the decision-making process. The PyTorch Forecasting package [9] has on open-source implementation of the model. It comes with its own optimizer, selecting the optimal number of attention heads, the network’s hidden size, and learning and dropout rates.
@@ 4 Search strategies
Classical approaches to hyperparameter tuning in machine learning, such as grid search and random search, have been widely used due to their simplicity [10]. However, these methods can be computationally expensive and inefficient, especially when dealing with high-dimensional search spaces or expensive objective functions [11]. Many approaches have been introduced to improve performance, including but not limited to: particle swarm optimization [12], using a simple neural predictor [13], NAS without training [14], Bayesian optimization, reinforcement learning, and the hyperband method. An extensive survey on over 1000 NAS papers [2] favorably featured the three latter approaches among the strategies for chain-structured search spaces, and these are the methods we explored in our work. The survey also prioritized cell-based, hierarchical search spaces and one-shot techniques, but we believe chain-structured spaces make more sense for small datasets like time series.
@@ 4.1 Bayesian optimization
Bayesian optimization has emerged as a popular alternative to grid and random search for hyperparameter tuning, as it efficiently explores the search space and intelligently guides the optimization process [15]. One widely used technique in Bayesian optimization is based on Gaussian processes (GPs), which fit the objective function using Gaussian processes as probabilistic models and leverages acquisition functions to balance exploration and exploitation [16]. GPs provide a flexible, nonparametric tool for modeling complex functions, allowing for uncertainty quantification and adaptation to new data [17].
However, GP-based methods have some limitations, such as the inability to handle categorical features or dependent parameters without ad hoc modifications [18]. This has led to the development of alternative methods, such as the Tree-structured Parzen Estimator (TPE), which has gained attention for its scalability and effectiveness [11]. TPE models the search space using hierarchical Parzen estimators, which adaptively partition the space to model complex, high-dimensional functions [11]. A simplified high-level version of the TPE algorithm for neural networks optimization is presented in Algorithm 1. Line 6 in Algorithm 1 makes intuitive sense as ‘choose a point where $x_{next}$ is most likely “good” and least likely “bad”’. Theoretically, it is justified since optimizing the criterion of Expected Improvement (EI) – the goal of TPE – is equivalent to maximizing $l(x)/g(x)$ at each step of the algorithm [11]. See [11] for a more detailed introduction to TPE and [19] for an in-depth tutorial.
Algorithm 1 Simplified TPE NAS
0:  A search space $S$ of neural network hyperparameters, number of trials $n_{trials}$, number of initial random samples $n_{init}$
0:  Results $R$ for each hyperparameter configuration tested
1:  Randomly select $n_{init}$ configurations from $S$.
2:  Train and test neural nets with the selected configurations, recording results in $R$
3:  for $n=1$ to $n_{trials}$ do
4:     Split $R$ into $R_{good}$ and $R_{bad}$ based on their performance according to the test metric.
5:     Fit two density models: $l(x)$ for $R_{good}$ and $g(x)$ for $R_{bad}$
6:     Choose the next configuration $x_{next}$ by maximizing $l(x_{next})/g(x_{next})$
7:     Train and test a neural net with the new configuration, updating $R$
8:  end for
9:  return  $R$
TPE is particularly suitable for addressing the shortcomings of GP-based methods, as it can naturally handle categorical features and dependent parameters without requiring extensive adaptations. Performance comparisons between TPE, GP-based methods, and classical search techniques have shown that TPE often yields improved results across various tasks and datasets [20].
A disadvantage of Bayesian optimization in general and TPE especially is the high theoretical complexity of the method when compared with simpler techniques, often for not much gain [2, 13]. Thankfully, a good well-supported practical implementation of TPE (that we used for our comparison) is available in the Optuna package, a versatile optimization library that has demonstrated its utility in various machine learning tasks [21].